OpenStudy (anonymous):

The equation of a line in slope-intercept form that passes through (6,1) and (-10,9)

OpenStudy (aroub):

9-1/-10-6 8/-16 1/-2

OpenStudy (anonymous):

y2-y1/x2-x1 = Slope formula y2 = 9 y1 = 1 x2 = -10 x1 = 6 So we have: 9-1/-10-6 x=3/11 And the equation of your line is: y = 3/11x + b, where b is the y-intercept. http://www.tutorsean.net

OpenStudy (anonymous):

How did you get 3/11?

OpenStudy (anonymous):

3/11 is the slope and he got it with this formula M=y2-y1/x2-x1 x1 and y1 being (6,1) and x2 and y2 being (-10,9)

OpenStudy (anonymous):

would the slope not be 8/-16 like aroub said?

OpenStudy (aroub):

oh my answer is incorrect?!?!

OpenStudy (aroub):

hey mo429 i checked my answer i guess its totally correct!! :)

OpenStudy (anonymous):

Haha thank you! redoing an exam in two days, CRAMMING

OpenStudy (anonymous):

No mo429 is correct. I made a dyslexia mistake in my head when I was figuring it. I transposed the numbers and mixed up the x and y when I worked it. Sorry about that.